A sample of nitrogen gas, N2, is collected in a 100 mL container at a pressure of 688 mm Hg and a temperature of 565 °C. How many grams of nitrogen gas are present in this sample? Round-off your answer up to 3 significant figures





1
Expert's answer
2021-12-14T11:40:01-0500

According to the ideal gas law:

"PV=nRT=mRT\/M",

where R = 8.314 J/(mol*K), the universal gas constant.


We should change all the given data to the SI units:

P = 688 mm Hg = 688 x 133.322 = 91725.536 Pa (as 1 mm Hg = 133.322 Pa)

V = 100 mL = 100 x 10-6 = 10-4 m3 (as 1 mL = 10-6 m3)

T = 565 °C = 565 + 273.15 = 838.15 K

M(N2) = 28.02 g/mol


"m = PVM\/RT =" 91725.536 x 10-4 x 28.02/(8.314 x 838.15) = 0.037 g